# Dense Conjugacy Classes and Stability of Locally Finite Graphs

### v1

@@ 4.3 Wedge decomposition
In this section, we prove the existence and uniqueness of a natural wedge decomposition for graphs whose maximal ends are all stable. A version of this decomposition for the end space of a surface was done by Schaffer-Cohen in [SC, Lemma 4.5] (and [BV]). We start by showing that locally finite graphs with stable maximal ends have finitely many maximal ends.
@@ 4.4 Ordered signatures
Figure 9: The graph $(\{Y_{n}\}_{n},a)\to\mathcal{C}$, which can also be written as $((1\vee C)\to 1)\vee((1\vee C)\to 1)$. In the latter signature, the self-similar locally finite graphs 11, $C$, and $(1\vee C)\to 1$, whose end spaces are the three local structures in the depicted graph, all appear as wedge components.
@@ 5 Dense conjugacy classes in $\text{Maps}(X)$
We now turn our attention to the question of which locally finite graphs $X$ are such that $\text{Maps}(X)$ has a dense conjugacy class. As mentioned in the introduction, there exists a complete classification in the surface setting due to [LV] and [HHMRSV].
@@
The JEP is useful because if $G$ is a Polish group, as in the case of $\text{Maps}(X)$ by Proposition 2.7, then $G$ has a dense conjugacy class if and only if $G$ has the JEP [KR, Theorem 2.1]. The topology of $\text{Maps}(X)$ is much more abstract than the topology of the graph $X$ itself, and when we will use the JEP to show that $\text{Maps}(X)$ does not have a dense conjugacy class, the sets $U$ and $V$ are going to be constructed directly using the topology on $X$. Thus, the JEP serves as a dictionary between the topology of $\text{Maps}(X)$ and the topology of $X$.
@@
Theorem 5.1 implies that the mapping class group of a surface with a unique maximal end, with genus zero or infinity, and which has self-similar end space, has a dense conjugacy class. The proof in [LV] of this statement is in fact the same in our setting. We give a sketch of their argument below.
@@
[LV] technically shows this result in the setting where the end space is self-similar, howeaver, this is true in our setting by Proposition 4.4. We may assume by Proposition 2.11 that $X$ is in standard form. Let $\mu$ be the maximal end of $X$. Given any edge $e$ in the underlying tree of $X$, the complement in $X$ of the interior of $e$ consists of two connected components; let $\Omega_{e}$ be the component such that $\mu\in\partial\Omega_{e}$, and let $\Sigma_{e}$ be the other one. Let $G$ be the subgroup of $\text{Maps}(X)$ consisting of elements with representatives restricting to the identity on some $\Omega_{e}$ for some edge $e$ in the underlying tree of $X$. The proof begins by proving the following claim.
@@
By [LV], $S$ must either be a two sphere or have genus either zero or infinite, a unique maximal end, and a self-similar end space. If $S$ is a 2-sphere, then $X$ is proper homotopy equivalent to a point, and hence has trivial mapping class group. Otherwise, $X$ must satisfy the conditions of Proposition 5.4, in which case $\text{Maps}(X)$ also has a dense conjugacy class. ∎
@@
Because $C$ is a tree, by Proposition 2.14, it suffices to show that the homeomorphisms of Cantor space has a dense conjugacy class. This was proven in [GW, Theorem 2.6]. ∎
@@ 5.3 Flux maps
In this section, we prove that a large class of mapping class groups do not contain a dense conjugacy class, proving Theorem 1.2 (1). Our strategy is to find a non-trivial continuous group homomorphism from $\text{Maps}(X)$ to $\mathbb{Z}$, which we call a flux map, similar to the flux maps defined in [DHK, Section 7]. The kernel of the flux map is a proper open normal subgroup of $\text{Maps}(X)$, and, by Lemma 5.3, this implies that $\text{Maps}(X)$ has no dense conjugacy class.
@@
As a first example, consider the graph $o(1)\vee o(1)$, which is a line with a copy of $S^{1}$ attached to each vertex. A loop shift, formally defined in [DHK, Section 3.4], is a proper homotopy equivalence of $o(1)\vee o(1)$ which sends each copy of $S^{1}$ to the adjacent one as in Figure 13, and it’s image under the flux map that counts loops is one. On the other hand, a similar map defined on the graph $o(\omega+1)\vee o(\omega+1)$ sending non-maximal ends to adjacent ones as in Figure 14 has infinitely many loops passing between the two maximal ends, and thus, counting loops would not yield a map to $\mathbb{Z}$. Instead, we count ends in $E(\partial o(1))$. Near the beginning of Section 5, we mentioned that the graph $o(1)$ can be thought of as a ray dominating copies of $S^{1}$ as motivation for considering immersed loops to be minimal ends, and this setting provides more motivation. Note that in both examples, the ends we count should satisfy a maximality condition to avoid the possibility of infinite ends or loops shifting towards an end. This becomes more apparent if one replaces the copies of $S^{1}$ in Figure 14 by rays. In light of this, we define a gcd, or greatest common divisor of a pair of ends.
@@
For example, we have that $R_{1}\in\gcd(o(1),o(1))$ in $o(1)\vee o(1)$ (see Figure 13), and while $R_{1}$ is not a gcd of the two maximal ends in Figure 14, $\partial o(1)$ is. Additionally, $\partial 1$ is a gcd of the two maximal end types in both graphs in Figure 15. Note that gcds need not be unique. In fact, the two maximal ends of the graph $(\{\bigvee_{i=0}^{n}X_{i}\}_{n}\to 1)\vee(\{\bigvee_{i=0}^{n}X_{i}\}_{n}\to 1)$, where the $X_{i}$ are as constructed in Lemma 4.14, have countably infinitely many gcds. Moreover, the two maximal ends of $((\{X_{n}\}_{n},a)\to G)\to(1\vee 1)$, where the $X_{n}$ are as constructed in Lemma 4.14, $a$ is a bijective attaching map, and $G$ is a Cantor tree, have uncountably many gcds.
@@
Constructing $T$ and $Z_{\nu}$: We define a subgraph $T$ of $X$ which contains no ends in $E(\lambda)$, similar to how an underlying tree consists of $X$ without loops. Because $\lambda$ is not of Cantor type, it follows from Lemma 3.6 that for each end $\nu\in E(\lambda)$, there exists a clopen neighborhood $U_{\nu}$ of $\nu$ such that $U_{\nu}\cap E(\lambda)=\{\nu\}$, i.e., such that $U_{\nu}$ is disjoint from all other ends in $E(\lambda)$. The neighborhoods $U_{\nu}$ can be made arbitrarily small, and thus we may assume that the $U_{\nu}$ are pairwise disjoint. Consider the clopen cover $\{U_{\nu}\}_{\nu\in E(\lambda)}$ of $E(\lambda)$. The complement of this cover in the end space, given by $\partial X\setminus(\bigcup_{\nu\in E(\lambda)}U_{\nu})$, is closed in $\partial X$. Therefore, there exists a subgraph $T$ of $X$ whose end space is $\partial X\setminus(\bigcup_{\nu\in E(\lambda)}U_{\nu})$. Intuitively, the graph $T$ is obtained from $X$ by removing all ends in $E(\lambda)$. See Figure 15 for examples when $X\cong 1\to(1\vee C)$ and $(\omega^{2}+1)\vee((1\to C)\to o(1))$. Define $Z_{\nu}$ to be a graph with end space equal to $U_{\nu}$ for each $\nu\in E(\lambda)$. Assume, by potentially altering $X$ up to proper homotopy equivalence, that $X$ is a wedge product of $T$ with each $Z_{\nu}$, and for each $\nu\in E(\lambda)$, let $x_{\nu}$ be vertex in $T$ which is the wedge point connecting $Z_{\nu}$ to the rest of $X$.
@@
    For all ends $\lambda^{\prime}\in\partial Z_{\nu}$, $\lambda^{\prime}\not\succ\lambda$, because any neighborhood of an end dominating $\lambda$ contains infinitely many points in $E(\lambda)$.
@@
The flux maps in [DHK, Section 7] are defined using a tool called corank, which calculates how many loops are in one subgraph containing $\mu_{1}$ but not another. We define corank in our setting in a similar vein.
@@
is finite because then the pair $(m,n)$ would be admissible relative to $f$, where $m=\max\{\lceil p\rceil,n\}$. This is because $\partial X_{m}\cap E(\lambda)$ must contain both $\partial f(X_{n})\cap E(\lambda)$ and $\partial X_{n}\cap E(\lambda)$ by construction. Suppose for contradiction that $p$ is not finite. Then the full subgraph generated by all of the $x_{\nu}\in\mathcal{V}(T\cap f(X_{n})\cap Y_{2})$ (where $\nu\in E(\lambda)$) is infinite. Thus, by Kőnig’s lemma [K], that subgraph contains an end $\lambda^{\prime}$. By Remark 5.19, both $\partial Y_{2}$ and $\partial f(X_{n})$ are clopen in $\partial X$, and so $\lambda^{\prime}\in\partial Y_{2}$ and $\lambda^{\prime}\in\partial f(X_{n})$. Also, $\lambda\prec\lambda^{\prime}$, since a sequence of vertices $x_{\nu}$ converges to $\lambda^{\prime}$. By Lemma 5.9, the intersection $E(\lambda^{\prime})\cap\partial X_{n}$ is nonempty. But then by Remark 5.19, $\partial Y_{1}\cap E(\lambda^{\prime})\neq\emptyset$, contradicting the assumption that $\lambda$ is a gcd. Thus the pair $(\lceil p\rceil,n)$ is admissible. ∎
@@
The proof of [DHK, Lemma 7.11] is sufficient for this claim, and we summarize it here. First, the quantity $\text{cork}(X_{m},X_{n})-\text{cork}(X_{m},f(X_{n}))$ is indeed an element of $\mathbb{Z}$ by Claim 5.24, and it is immediate that $\phi_{m,n}(id)=0$. For $f,g\in\text{PHE}(X)$ and $n\in\mathbb{Z}$, there exists $m\in\mathbb{Z}$ such that $(m,n)$ is admissible relative to all three maps $f$, $g$, and $fg$ by Claim 5.23. The rest of the proof consists of algebraic manipulations and making use of the additivity of corank mentioned in Example 5.21. Lastly, the continuity of $\phi_{m,n}$ follows as any homeomorphism to $\mathbb{Z}$ is continuous [D, Theorem 1]. ∎
@@
The first statement follows from the fact that if $f$ and $g$ are properly homotopic, then they induce the same homeomorphism on $\partial X$. The moreover statement follows from the same proof as in [DHK, Lemma 7.10] which we briefly summarize. Additivity of corank (see Example 5.21) gives that $\phi_{m^{\prime},n}(f)=\phi_{m,n}(f)$ for any $f$ through algebraic manipulations similar to those in triangle inequality proofs, and thus it suffices to show that $\phi_{m,n}=\phi_{m,n^{\prime}}$. This follows from additivity and invariance after applying proper homotopy equivalences (see Example 5.21) through similar manipulations. ∎
@@
Because $\Phi$ is a continuous group homomorphism, the kernel of $\Phi$ is a proper open normal subgroup of $\text{Maps}(X)$. By Lemma 5.3, $\text{Maps}(X)$ has no dense conjugacy class. ∎
@@
Each $X_{n}$ corresponds to a subgroup $A_{n}:=\pi_{1}(X_{n},x_{0})$ of $\pi_{1}(X,x_{0})$. Moreover, this subgroup is a free factor of $\pi_{1}(X,x_{0})$, namely, there exist subgroups $H_{n}<\pi_{1}(X,x_{0})$ such that $\pi_{1}(X,x_{0})\cong A_{n}*H_{n}$ for each $n$. Furthermore, for $n,m\in\mathbb{Z}$ with $n<m$, $A_{n}\leq A_{m}$, and $A_{n}$ is a free factor of $A_{m}$ [DHK, Lemma 7.2]. Let
@@
Corank is additive and invariant after applying proper homotopy equivalences. A pair of integers $(m,n)$ with $n<m$ is admissible relative to a proper homotopy equivalence $f$ if $A_{n}$ and $f_{*}(A_{n})$ are free factors of $A_{m}$. The proofs of Claims 5.23, 5.24, 5.25, and 5.26 are the same after making the appropriate changes from ends to loops. ∎
@@
If $\mu_{1}\not\sim\mu_{2}$, then the result follows from Propositions 5.18 and 5.28. Otherwise, suppose that $\mu_{1}\sim\mu_{2}$. If $E(\mu_{1})$ is finite, then $|E(\mu_{1})|>1$, so Theorem 5.10 implies the result.
@@
    Consider the graph with wedge decomposition $(X\to 1)\vee(X\to o(1))$, where $X$ is a locally finite tree with a unique maximal end which is not self similar (and not stable) such as the one constructed in [MRshort]. A gcd for the two maximal ends is the maximal end in any copy of $X$. In particular, the gcd need not be a stable end type.
@@
This yields the following generalization of Theorem 1.2 (1).
@@
Let $X$ be a locally finite graph, and suppose that $A_{1},A_{2}\subset\partial X$, $\lambda\in\partial X$, and $H\leq\text{Maps}(X)$ satisfy the three conditions above. Then $H$ does not have a dense conjugacy class.
@@
By the same methods as in Propositions 5.18 and 5.28, the map $\Phi$ is a continuous group homomorphism, and so the kernel of $\Phi$ is a proper open normal subgroup of $H$. By Lemma 5.3, $H$ has no dense conjugacy class. ∎
Theorem 5.32 implies Propositions 5.18 and 5.28 by letting $\nu_{1}$ and $\nu_{2}$ be maximal ends of different types and setting $A_{i}=E(\nu_{i})$ for $i=1,2$. Moreover, we have the following corollary.

### v2

@@ 4.3 Wedge decomposition
In this section, we prove the existence and uniqueness of a natural wedge decomposition for graphs whose maximal ends are all stable. A version of this decomposition for the end space of a surface was done by Schaffer-Cohen in [SCH24, Lemma 4.5] (and [BV23]). We start by showing that locally finite graphs with stable maximal ends have finitely many maximal ends.
@@ 4.4 Ordered signatures
Figure 9: The graph $(\{Y_{n}\}_{n},a)\to\mathcal{C}$, which can also be written as $((1\vee C)\to 1)\vee((1\vee C)\to 1)$. In the latter signature, the self-similar locally finite graphs $1$, $C$, and $(1\vee C)\to 1$, whose end spaces are the three local structures in the depicted graph, all appear as wedge components.
@@ 5 Dense conjugacy classes in $\text{Maps}(X)$
We now turn our attention to the question of which locally finite graphs $X$ are such that $\text{Maps}(X)$ has a dense conjugacy class. As mentioned in the introduction, there exists a complete classification in the surface setting due to [LV22] and [HHM+22].
@@
The JEP is useful because if $G$ is a Polish group, as in the case of $\text{Maps}(X)$ by Proposition 2.7, then $G$ has a dense conjugacy class if and only if $G$ has the JEP [KR07, Theorem 2.1]. The topology of $\text{Maps}(X)$ is much more abstract than the topology of the graph $X$ itself, and when we will use the JEP to show that $\text{Maps}(X)$ does not have a dense conjugacy class, the sets $U$ and $V$ are going to be constructed directly using the topology on $X$. Thus, the JEP serves as a dictionary between the topology of $\text{Maps}(X)$ and the topology of $X$.
@@
Theorem 5.1 implies that the mapping class group of a surface with a unique maximal end, with genus zero or infinity, and which has self-similar end space, has a dense conjugacy class. The proof in [LV22] of this statement is in fact the same in our setting. We give a sketch of their argument below.
@@
[LV22] technically shows this result in the setting where the end space is self-similar, howeaver, this is true in our setting by Proposition 4.4. We may assume by Proposition 2.11 that $X$ is in standard form. Let $\mu$ be the maximal end of $X$. Given any edge $e$ in the underlying tree of $X$, the complement in $X$ of the interior of $e$ consists of two connected components; let $\Omega_{e}$ be the component such that $\mu\in\partial\Omega_{e}$, and let $\Sigma_{e}$ be the other one. Let $G$ be the subgroup of $\text{Maps}(X)$ consisting of elements with representatives restricting to the identity on some $\Omega_{e}$ for some edge $e$ in the underlying tree of $X$. The proof begins by proving the following claim.
@@
By [LV22], $S$ must either be a two sphere or have genus either zero or infinite, a unique maximal end, and a self-similar end space. If $S$ is a 2-sphere, then $X$ is proper homotopy equivalent to a point, and hence has trivial mapping class group. Otherwise, $X$ must satisfy the conditions of Proposition 5.4, in which case $\text{Maps}(X)$ also has a dense conjugacy class. ∎
@@
Because $C$ is a tree, by Proposition 2.14, it suffices to show that the homeomorphisms of Cantor space has a dense conjugacy class. This was proven in [GW01, Theorem 2.6]. ∎
@@ 5.3 Flux maps
In this section, we prove that a large class of mapping class groups do not contain a dense conjugacy class, proving Theorem 1.2 (1). Our strategy is to find a non-trivial continuous group homomorphism from $\text{Maps}(X)$ to $\mathbb{Z}$, which we call a flux map, similar to the flux maps defined in [DHK23, Section 7]. The kernel of the flux map is a proper open normal subgroup of $\text{Maps}(X)$, and, by Lemma 5.3, this implies that $\text{Maps}(X)$ has no dense conjugacy class.
@@
As a first example, consider the graph $o(1)\vee o(1)$, which is a line with a copy of $S^{1}$ attached to each vertex. A loop shift, formally defined in [DHK23, Section 3.4], is a proper homotopy equivalence of $o(1)\vee o(1)$ which sends each copy of $S^{1}$ to the adjacent one as in Figure 13, and it’s image under the flux map that counts loops is one. On the other hand, a similar map defined on the graph $o(\omega+1)\vee o(\omega+1)$ sending non-maximal ends to adjacent ones as in Figure 14 has infinitely many loops passing between the two maximal ends, and thus, counting loops would not yield a map to $\mathbb{Z}$. Instead, we count ends in $E(\partial o(1))$. Near the beginning of Section 5, we mentioned that the graph $o(1)$ can be thought of as a ray dominating copies of $S^{1}$ as motivation for considering immersed loops to be minimal ends, and this setting provides more motivation. Note that in both examples, the ends we count should satisfy a maximality condition to avoid the possibility of infinite ends or loops shifting towards an end. This becomes more apparent if one replaces the copies of $S^{1}$ in Figure 14 by rays. In light of this, we define a gcd, or greatest common divisor of a pair of ends.
@@
For example, we have that $R_{1}\in\gcd(o(1),o(1))$ in $o(1)\vee o(1)$ (see Figure 13), and while $R_{1}$ is not a gcd of the two maximal ends in Figure 14, $\partial o(1)$ is. Additionally, $\partial 1$ is a gcd of the two maximal end types in both graphs in Figure 15. Note that gcds need not be unique. In fact, the two maximal ends of the graph $(\{\bigvee_{i=0}^{n}X_{i}\}_{n}\to 1)\vee(\{\bigvee_{i=0}^{n}X_{i}\}_{n}\to 1)$, where the $X_{i}$ are as constructed in Lemma 4.14, have countably infinitely many gcds.
@@
Constructing $T$ and $Z_{\nu}$: We define a subgraph $T$ of $X$ which contains no ends in $E(\lambda)$, similar to how an underlying tree consists of $X$ without loops. Because $\lambda$ is not of Cantor type, it follows from Lemma 3.6 that any stable neighborhood $U_{\nu}$ of each end $\nu\in E(\lambda)$ is such that $U_{\nu}\cap E(\lambda)=\{\nu\}$, i.e., such that $U_{\nu}$ is disjoint from all other ends in $E(\lambda)$. The neighborhoods $U_{\nu}$ can be made arbitrarily small, and thus we may assume that the $U_{\nu}$ are pairwise disjoint. Consider the clopen cover $\{U_{\nu}\}_{\nu\in E(\lambda)}$ of $E(\lambda)$. The complement of this cover in the end space, given by $\partial X\setminus(\bigcup_{\nu\in E(\lambda)}U_{\nu})$, is closed in $\partial X$. Therefore, there exists a subgraph $T$ of $X$ whose end space is $\partial X\setminus(\bigcup_{\nu\in E(\lambda)}U_{\nu})$. Intuitively, the graph $T$ is obtained from $X$ by removing all ends in $E(\lambda)$. See Figure 15 for examples when $X\cong 1\to(1\vee C)$ and $(\omega^{2}+1)\vee((1\to C)\to o(1))$. Define $Z_{\nu}$ to be a graph with end space equal to $U_{\nu}$ for each $\nu\in E(\lambda)$. Note that for each $\nu\in E(\lambda)$, $\nu\sim\partial Z_{\nu}$. Assume, by potentially altering $X$ up to proper homotopy equivalence, that $X$ is a wedge product of $T$ with each $Z_{\nu}$, and for each $\nu\in E(\lambda)$, let $x_{\nu}$ be vertex in $T$ which is the wedge point connecting $Z_{\nu}$ to the rest of $X$.
@@
    For all ends $\lambda^{\prime}\in\partial Z_{\nu}$, $\lambda^{\prime}\not\succ\lambda$ by Proposition 4.4.
@@
The flux maps in [DHK23, Section 7] are defined using a tool called corank, which calculates how many loops are in one subgraph containing $\mu_{1}$ but not another. We define corank in our setting in a similar vein.
@@
is finite because then the pair $(m,n)$ would be admissible relative to $f$, where $m=\max\{\lceil p\rceil,n\}$. This is because $\partial X_{m}\cap E(\lambda)$ must contain both $\partial f(X_{n})\cap E(\lambda)$ and $\partial X_{n}\cap E(\lambda)$ by construction. Suppose for contradiction that $p$ is not finite. Then the full subgraph generated by all of the $x_{\nu}\in\mathcal{V}(T\cap f(X_{n})\cap Y_{2})$ (where $\nu\in E(\lambda)$) is infinite. Thus, by Kőnig’s lemma [KÖN27], that subgraph contains an end $\lambda^{\prime}$. By Remark 5.19, both $\partial Y_{2}$ and $\partial f(X_{n})$ are clopen in $\partial X$, and so $\lambda^{\prime}\in\partial Y_{2}$ and $\lambda^{\prime}\in\partial f(X_{n})$. Also, $\lambda\prec\lambda^{\prime}$, since a sequence of vertices $x_{\nu}$ converges to $\lambda^{\prime}$. By Lemma 5.9, the intersection $E(\lambda^{\prime})\cap\partial X_{n}$ is nonempty. But then by Remark 5.19, $\partial Y_{1}\cap E(\lambda^{\prime})\neq\emptyset$, contradicting the assumption that $\lambda$ is a gcd. Thus the pair $(\lceil p\rceil,n)$ is admissible. ∎
@@
The proof of [DHK23, Lemma 7.11] is sufficient for this claim, and we summarize it here. First, the quantity $\text{cork}(X_{m},X_{n})-\text{cork}(X_{m},f(X_{n}))$ is indeed an element of $\mathbb{Z}$ by Claim 5.24, and it is immediate that $\phi_{m,n}(id)=0$. For $f,g\in\text{PHE}(X)$ and $n\in\mathbb{Z}$, there exists $m\in\mathbb{Z}$ such that $(m,n)$ is admissible relative to all three maps $f$, $g$, and $fg$ by Claim 5.23. The rest of the proof consists of algebraic manipulations and making use of the additivity of corank mentioned in Example 5.21. Lastly, the continuity of $\phi_{m,n}$ follows as any homeomorphism to $\mathbb{Z}$ is continuous [DUD61, Theorem 1]. ∎
@@
The first statement follows from the fact that if $f$ and $g$ are properly homotopic, then they induce the same homeomorphism on $\partial X$. The moreover statement follows from the same proof as in [DHK23, Lemma 7.10] which we briefly summarize. Additivity of corank (see Example 5.21) gives that $\phi_{m^{\prime},n}(f)=\phi_{m,n}(f)$ for any $f$ through algebraic manipulations similar to those in triangle inequality proofs, and thus it suffices to show that $\phi_{m,n}=\phi_{m,n^{\prime}}$. This follows from additivity and invariance after applying proper homotopy equivalences (see Example 5.21) through similar manipulations. ∎
@@
Claim 5.28.
The flux map $\Phi\colon\text{Maps}(X)\to\mathbb{Z}$ is non-trivial.
Proof.
An element of $\text{Maps}(X)$ that is mapped non-trivially to $\mathbb{Z}$ under $\Phi$ can be constructed using [MR23, Observation 4.9] on the collection $\{U_{\nu}\}_{\nu\in E(\lambda)}$. ∎
@@
Because $\Phi$ is a non-trivial continuous group homomorphism, the kernel of $\Phi$ is a proper open normal subgroup of $\text{Maps}(X)$. By Lemma 5.3, $\text{Maps}(X)$ has no dense conjugacy class. ∎
@@
Each $X_{n}$ corresponds to a subgroup $A_{n}:=\pi_{1}(X_{n},x_{0})$ of $\pi_{1}(X,x_{0})$. Moreover, this subgroup is a free factor of $\pi_{1}(X,x_{0})$, namely, there exist subgroups $H_{n}<\pi_{1}(X,x_{0})$ such that $\pi_{1}(X,x_{0})\cong A_{n}*H_{n}$ for each $n$. Furthermore, for $n,m\in\mathbb{Z}$ with $n<m$, $A_{n}\leq A_{m}$, and $A_{n}$ is a free factor of $A_{m}$ [DHK23, Lemma 7.2]. Let
@@
Corank is additive and invariant after applying proper homotopy equivalences. A pair of integers $(m,n)$ with $n<m$ is admissible relative to a proper homotopy equivalence $f$ if $A_{n}$ and $f_{*}(A_{n})$ are free factors of $A_{m}$. The proofs of Claims 5.23, 5.24, 5.25, and 5.26 are the same after making the appropriate changes from ends to loops, and Claim 5.28 follows from the existence of loop shifts [DHK23, Section 3.4]. ∎
@@
If $\mu_{1}\not\sim\mu_{2}$, then the result follows from Propositions 5.18 and 5.29. Otherwise, suppose that $\mu_{1}\sim\mu_{2}$. If $E(\mu_{1})$ is finite, then $|E(\mu_{1})|>1$, so Theorem 5.10 implies the result.
@@
    Consider the graph with wedge decomposition $(X\to 1)\vee(X\to o(1))$, where $X$ is a locally finite tree with a unique maximal end which is not self-similar (and not stable) such as the one constructed in [MR24]. Because the maximal end in any copy of $X$ is not stable, ends of this type are not technically gcds. The stability assumption is only used in Claim 5.28, and by shifting copies of $X$ in this graph, it is clear that there still exists non-trivial flux maps. See Remark 5.33 for the general principle.
@@
Remark 5.33.
The flux map $\Phi$ is non-trivial if there exists a clopen, pairwise disjoint, and pairwise homeomorphic cover $\{U_{\nu}\}_{\nu\in E(\lambda)}$ of $E(\lambda)$ in $\partial X$ such that $U_{\nu}\cap E(\lambda)$ is a singleton for all $\nu$. Such a cover always exists if $\lambda$ is stable.
The following is a generalization of Theorem 1.2 (1).
@@
Let $X$ be a locally finite graph, and suppose that $A_{1},A_{2}\subset\partial X$, $\lambda\in\partial X$, and $H\leq\text{Maps}(X)$ satisfy the three conditions above and the conditions of Remark 5.33. Then $H$ does not have a dense conjugacy class.
@@
By the same methods as in Propositions 5.18 and 5.29, the map $\Phi$ is a continuous group homomorphism, and so the kernel of $\Phi$ is a proper open normal subgroup of $H$. By Lemma 5.3, $H$ has no dense conjugacy class. ∎
Theorem 5.34 implies Propositions 5.18 and 5.29 by letting $\nu_{1}$ and $\nu_{2}$ be maximal ends of different types and setting $A_{i}=E(\nu_{i})$ for $i=1,2$. Moreover, we have the following corollary.